A metal was irradiated by light of frequency 3.2 X 10^15 s^-1. The photoelectron produced has its KR, 2 times the KE of photoelectron which was produced when the same metal was irradiated with a light of frequency 2.0 X 10^15 s^-1. What is its work function?

Grade:11

1 Answers

Ramesh V
14 years ago

K.E = hv -W

w is work fucnction in ev

h is planks constant = 4.14x10-15 ev-sec

K.E1 = hv1 -W  and K.E2 = hv2 -W

(hv1 -W) = 2.(hv2 -W )

W = h(2v2 -  v1)

     = 4.14x10-15 * (4 - 3.2)*x1015

     Work function = 3.312 eV

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch